OpenStudy (anonymous):

how do you convert ln(1/e)=-1 to exponential form

OpenStudy (mertsj):

What is the base?

OpenStudy (anonymous):

i dont see any base

OpenStudy (anonymous):

you raise both sides to the exponential e and you get e ^ [ln (1/e)] = e ^ (-1) Having this, you should know that e and ln are inverses of each other and you just get 1/e = e ^ (-1)

OpenStudy (precal):



OpenStudy (precal):